# Models That Prove Their Own Correctness

### v3

@@ 3.2 Self-Proving models
We define the Verifiability of a model $P_{\theta}$ with respect to an input distribution $\mu$ and a verifier $V$. Intuitively, Verifiability captures the ability of the model to prove the correctness of its answer $y\sim P_{\theta}(x)$, when the input $x$ is sampled from $\mu$. We refer to models capable of proving their own correctness as Self-Proving models. Notice that, as in Definition 3.2, the verifier is fixed and agnostic to the choice of the Self-Proving model.
@@
Fix a verifier $V$ for a ground-truth $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$ as in Definition 3.2, and a distribution $\mu$ over inputs $\Sigma^{\ast}$. The Verifiability of a model $P_{\theta}\colon\Sigma^{\ast}\to\Sigma^{\ast}$ is defined as
@@
Notice that the ground-truth $F^{\ast}$ does not appear in Definition 3.4 except for the first sentence. Indeed, once it is established that $V$ is a verifier for $F^{\ast}$ (as per Definition 3.2), then Verifiability w.r.t $V$ implies correctness w.r.t $F^{\ast}$: Consider any input distribution $\mu$, ground-truth $F^{\ast}$, and a verifier $V$ for $F^{\ast}$ with soundness error $s$. By a union bound, if model $P_{\theta}$ is $\beta$-Verifiable, then it is $(\beta-s)$-correct. That is to say, Verifiability is formally a stronger guarantee than correctness when $V$ has small soundness error $s$.
As depicted in Figure 1, a Self-Proving model $P_{\theta}$ plays a dual role: first, it generates an output $y\sim P_{\theta}(x)$, and then it proves the correctness of this output to $V$. Note also that Self-Provability is a feature of a model, unlike completeness and soundness which are features of a verifier (see Table 1).
@@
4 Learning Self-Proving Autoregressive Models
The challenge is to align the model with a verifier. We assume that the learner has access to input samples $x\sim\mu$ and correct outputs $F^{\ast}(x)$, as well as the verifier specification (code). Additionally, the learner can emulate the verifier, as the latter is computationally efficient (Remark 3.3).
Our focus is on autoregressive sequence-to-sequence (Self-Proving) models $P_{\theta}$. Such models generate their output by recursively prompting a randomized sampling from a base distribution $p_{\theta}$ over tokens $\Sigma$. For an input $z\in\Sigma^{\ast}$, the output $w\sim P_{\theta}(z)$ is generated as follows:
  • Sample $w_{1}\sim p_{\theta}(z)$.
  • Let $j=1$. While $w_{j}$ is not the end-of-sequence token $\mathtt{EOS}\in\Sigma$:
    • Sample $w_{j+1}\sim p_{\theta}(zw_{1}\cdots w_{j})$.
    • Update $j\coloneqq j+1$.
  • Output $w=w_{1}w_{2}\cdots w_{j}$.
For any $z\in\Sigma^{\ast}$, it is useful to consider the vector of log-probabilities over $\Sigma$, denoted by $\log p_{\theta}(z)\in\mathbb{R}^{|\Sigma|}$. We assume that each coordinate in this vector is differentiable with respect to $\theta$.
Our general approach is inspired by Reinforcement Learning from Human Feedback (Christiano et al., 2017), a method for aligning models with human preferences, which has recently been used to align sequence-to-sequence models (Ouyang et al., 2022). However, there are two important differences between humans and algorithmic verifiers: (1) Verifiers are efficient algorithms which may be emulated by the learner. This is unlike humans, whose preferences are costly to obtain. On the other hand, (2) verifiers make a single-bit decision at the end of an interaction, but cannot guide the prover (model) in intermediate rounds. In RL terms, this is known as the exploration problem for sparse reward signals (e.g. Ladosz et al. 2022).
Section 4.1 introduces Transcript Learning (TL), a learning algorithm that overcomes the exploration problem mentioned in the second point under the assumption that the learner has access to transcripts of interactions in which the verifier accepts. We prove convergence bounds for TL (Section B.1) and analyze it experimentally (Section 5).
Access to accepting transcripts is a reasonable assumption, for example, when there is an efficient honest prover that can generate such transcripts (Goldwasser et al., 2015). When there is no access to accepting transcripts, we propose Reinforcement Learning from Verifier Feedback (Section 4.2).
4.1 Transcript Learning
We present an algorithm for learning Self-Proving models which uses access to a distribution of accepting transcripts. This is a reasonable assumption to make when the honest prover $P^{\ast}$ (see Definition 3.2) is efficient, as in the case of public-coin Doubly-Efficient Interactive Proof systems as defined by Goldwasser et al. (2015) and developed in other theoretical (e.g. Goldreich and Rothblum 2018) and applied (e.g. Zhang et al. 2021) works. In this case, an honest prover $P^{\ast}$ can be run by the learner during training to collect accepting transcripts without incurring heavy computational cost. Alternatively, the learner may collect a dataset of accepting transcripts prior to learning (see Figure 4 in Appendix B).
The intuition behind Transcript Learning is that the interaction of the verifier and prover can be viewed as a sequence itself, which is called the transcript $\pi\in\Sigma^{\ast}$. The idea is to learn a model not just of $x\mapsto y^{\ast}$ for a correct output $y^{\ast}$, but of $x\mapsto y^{\ast}\pi^{\ast}$, where $\pi^{\ast}$ is a transcript of an interaction in which the verifier accepted.
In more detail, Transcript Learning (TL, Algorithm 1) requires access to an (honest) transcript generator $\mathcal{T}^{\ast}$. Given an input $x$, the generator $\mathcal{T}^{\ast}(x)$ samples a sequence $P^{\ast}\!(x)\pi^{\ast}\in\Sigma^{\ast}$ such that $\pi^{\ast}$ is an accepted transcript. TL trains a Self-Provable model by autoregressively optimizing towards generating accepting transcripts. At a very high level, it works by repeatedly sampling $x\sim\mu$ and transcript $y^{\ast}\pi^{\ast}\sim\mathcal{T}^{\ast}(x)$, and updating the logits $\log p_{\theta}$ towards agreeing with $y^{\ast}\pi^{\ast}$ via Gradient Ascent. We prove that, under certain conditions, it is expected to output a Self-Provable model.
Theorem 4.1 (Theorem B.5, informal).
Fix an input distribution $\mu$, a verifier $V$, a transcript generator $\mathcal{T}^{\ast}$, an autoregressive model family $\{P_{\theta}\}_{\theta}$ parameterized by $\theta\in\mathbb{R}^{d}$ for some $d\in\mathbb{N}$, and a norm |||||$\cdot$||| | ⋅ | | on $\mathbb{R}^{d}$. Assume that the agreement function $A\colon\mathbb{R}^{d}\to[0,1]$ defined by
$A(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ {\pi^{\ast}\sim\mathcal{T}^{\ast}(x)}\end{subarray}}[\mathrm{Transcript}(\left<V,P_{\theta}\right>(x))=\pi^{\ast}]$
is concave in $\theta$. For any $\varepsilon>0$, let $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ be upper-bounds such that the following conditions hold.
  • There exists $\theta^{\ast}\in\mathbb{R}^{d}$ with $||\theta^{\ast}||<B_{\mathrm{Norm}}$ such that $A(\theta^{\ast})\geq 1-\varepsilon/2$.
  • For all $\theta$, the logits of $P_{\theta}$ are $B_{\mathrm{Lip}}$-Lipschitz in $\theta$.
  • The total number of tokens sent by the prover to the verifier $V$ in any interaction is at most $C$.
Denote by $\bar{\theta}$ the output of Transcript Learning (Algorithm 1) running for $N$ interations, where
$N\geq 4\cdot C^{2}\cdot\frac{B_{\mathrm{Norm}}^{2}\cdot B_{\mathrm{Lip}}^{2}}{\varepsilon^{2}}$ (2)
and learning rate $\lambda=B_{\mathrm{Norm}}/CB_{\mathrm{Lip}}\sqrt{N}$. Then the expected Verifiability of $\bar{\theta}$ is at least $1-\varepsilon$.
The proof (Appendix B) goes by reduction to Stochastic Gradient Descent (SGD). We show (Lemma B.4) that the learner can use its only available tools—sampling honest transcripts, emulating the verifier, and differentiating the logits—to optimize the agreement $A(\theta)$. Specifically, this is done by accumulating gradients from the cross-entropy loss computed at each token. Since $A(\theta)$ lower bounds the Verifiability of $P_{\theta}$, the former can be used as a surrogate for the latter.
The conditions for Theorem 4.1 can be split into two. First, the standard conditions used to prove SGD convergence: convexity, [^4] Theorem 4.1 requires concavity because it guarantees maximization, rather than minimization. We leave it for future work to relax the differentiability assumption. $B_{\mathrm{Norm}}$-boundedness, and $B_{\mathrm{Lip}}$-Lipschitzness. Second, there is a bound $C$ on the communication complexity of the prover in the Interactive Proof system.
Quantitatively, the efficiency of TL is captured by the number of iterations $N$. It is desirable to minimize $N$, which is also the number of samples needed from the distribution $\mu$ and the transcript generator $\mathcal{T}^{\ast}$. The bound on $N$ in Equation 2 can be decomposed into the complexity of SGD ($B_{\mathrm{Norm}}^{2}B_{\mathrm{Lip}}^{2}/\varepsilon^{2}$), and communication complexity of the proof system $O(C^{2})$. Minimizing communication complexity has been an overarching goal in the study of proof systems (e.g. Goldreich and Håstad 1998; Goldreich et al. 2002; Reingold et al. 2021). Theorem 4.1 formally shows the benefit of communication-efficient proof systems in the context of Self-Proving models.
4.2 Reinforcement Learning from Verifier Feedback (RLVF)
As mentioned in Section 4.1, Transcript Learning uses access to an honest transcript generator to estimate gradients of (a lower bound on) the Verifiability of a model $P_{\theta}$.
Reinforcement Learning from Verifier Feedback (RLVF, Algorithm 2) estimates this gradient without access to a transcript generator. RLVF can be viewed as a modification of TL in which the learner emulates the interaction of the verifier with its own model $P_{\theta}$. Rather than directly sampling from the generator as in TL, it collects accepting transcripts by rejection sampling on emulated transcripts.
This rejection sampling means that RLVF requires its initial model $P_{\theta_{0}}$ to have Verifiability bounded away from 0, so that accepting transcripts are sampled with sufficient probability. Fortunately, such a Self-Proving base model can be learned using TL. This gives a learning paradigm in which a somewhat-Self-Proving base model is first learned with TL (with Verifiability $\delta>0$), and then “amplified” to a fully Self-Proving model using RLVF (cf. Nair et al. 2018).
We prove that RLVF learner can estimate the Verifiability gradient of $P_{\theta}$ using emulation alone in Lemma B.6. From a broader perspective, RLVF can be derived by viewing Self-Proving as a reinforcement learning problem in which the agent (prover) is rewarded when the verifier accepts. Indeed, RLVF is the Policy Gradient method (Sutton et al., 1999) for a verifier-induced reward. Convergence bounds for Policy Gradient methods are a challenging and active area of research (e.g. Agarwal et al. 2021), and so we leave the full analysis to future work.
4.3 Learning from annotated transcripts
However, when training Self-Proving model, it may be useful for it to first generate an “annotated” answer $\widetilde{a}$ which is then trimmed down to the actual answer $a$ to be sent to the verifier. We adapt Sections 3 and 4 to this setting in Appendix D, where we present Annotated Transcripts. The TL and RLVF algorithms naturally extend to annotated transcripts as well. Table 2 shows that annotations significantly improve performance of TL.
Annotations can be viewed as adding Chain-of-Thought (Wei et al., 2022). As a concrete example, consider our experiments on computing the GCD. As detailed in Section 5.2, a proof $\pi$ in this setting is the output of an iterative process—the extended Euclidean algorithm—starting from the input $x$: $x\mapsto\pi_{1}\mapsto\pi_{2}\mapsto\dots\mapsto\pi$. The annotation of the proof $\pi$ consists the first $T$ steps $(\pi_{1},\dots,\pi_{T})$ up to some fixed cutoff $T$. These are prepended to the proof and shown to the model during TL training. At inference time, the model is evaluated only on whether it generated the proof $\pi$ correctly.
5 Experimental Results
5.1 Setup: Training transformers to predict the GCD of two integers
Charton (2024) empirically studies the power and limitations of learning GCDs with transformers. We follow their setup and two conclusions on settings that make for faster learning: Training from the log-uniform distribution, and choosing a base of representation with many prime factors.
[…]
We choose the input distribution $\mu$ to be the log-uniform distribution on $[10^{4}]$, and train the transformer on sequences of the form $\mathbf{(x_{1}x_{2}y)}$, where $x_{1},x_{2}\sim\mu$ and $y=GCD(x_{1},x_{2})$. This is a scaling-down of Charton (2024), to allow single GPU training of Self-Proving transformers. In all of our experiments, we use a GPT model (Vaswani et al., 2017) with 6.3M parameters trained on a dataset of 1024K samples in batches of 1024. Full details are deferred to Appendix F.
Proving correctness of GCD.
Following Charton (2024) as a baseline, we find that transformers can correctly compute the GCD with over $99\%$ probability over $(x_{1},x_{2})\sim\mu$. To what extent can they prove their answer? To answer this question, we first devise a natural proof system based on Bézout’s theorem. Its specification and formal guarantees are deferred to Appendix E. We denote its verification algorithm by $V$, and highlight some important features of the experimental setup:
  • Completeness: For any $x_{1},x_{2},y\in[10^{4}]$ such that $y=GCD(x_{1},x_{2})$, there exists a proof $\mathbf{\pi}$ such that $V(\mathbf{x_{1}x_{2}y\pi})$ accepts. As detailed in Appendix E, the proof $\pi$ consists of a pair of integers who are Bézout coefficients for $x_{1},x_{2}$.
Table 2 shows our main experimental result, which has the following key takeaways:
  1. 1.
[…]
  2. 2.
  3. 3.
    Most efficient is Annotated Transcript Learning, which yielded a model with 96% Verifiability in 100K iterations.
We further investigate the effect of annotations next.
5.2 Models generalize beyond annotations
The proof $\mathbf{\pi}$ is annotated by including intermediate steps in its computation. Details are deferred to Appendix E; roughly speaking, we observe that the proof $\mathbf{\pi}$ for input $\mathbf{(a,b)}$ is obtained as the last element in a sequence $\mathbf{a,b,\pi_{1},\pi_{2},\dots}$ computed by the Euclidean algorithm. We annotate the proof $\mathbf{\pi}$ by prepending to it the sequence of Euclidean steps $\mathbf{(\pi_{1},\dots,\pi_{T})}$ up to some fixed cutoff $T$.
Figure 2 shows how $T$ affects the Verifiability of the learned model. As suggested by Lee et al. (2024), training the model on more intermediate steps results in better performance; in our case, increasing the number of intermediate steps $T$ yields better Self-Proving models. One might suspect that models only learn to execute the Euclidean algorithm in-context. To rule out this hypothesis, we derive an upper bound on the possible efficacy of such limited models. This bound is based on the Euclidean depth of integers $(x_{1},x_{2})$, which we define as the number of intermediate steps that the Euclidean algorithm makes before terminating on input $(x_{1},x_{2})$. Indeed, a model that only learns to compute (in-context) the simple arithmetic of the Euclidean algorithm would only be able to prove the correctness of inputs $(x_{1},x_{2})$ whose depth does not exceed the annotation cutoff $T$.
Figure 2 tells a different story: For each cutoff $T$, we estimate the probability that integers $x_{1},x_{2}\sim\mu$ have Euclidean depth at most $T$ on $10^{5}$ sampled pairs. Larger annotation cutoff $T$ increases Verifiability, but all models exceed their corresponding Euclidean depth bound.
5.3 Base of representation
As mentioned previously, Charton (2024) concludes that, for a given base of representation $B$, transformers correctly compute the GCD of integers $x_{1},x_{2}$ that are products of primes dividing $B$. Simply put, choosing a base $B$ with many different prime factors yields models with better correctness (accuracy), which suggests why base $B=210=2\cdot 3\cdot 5\cdot 7$ yielded the best results.
To test whether the factorization of $B$ has a similar effect on Verifiability as well, we train transformers on 68 bases varying the number of prime divisors $\omega(B)$ from $\omega(B)=1$ (i.e., $B$ is a prime power) to $\omega(B)=4$. Figure 3 shows that $\omega(B)$ correlates not just with correctness (Charton, 2024), but also with Verifiability. Although the finding is statistically significant (no overlapping error margins), the overall difference is by a few percentage points; we attribute this to the smaller (10%) number of samples on which models were trained, relative to our other experiments.
6 Limitations
Our experiments are focused on a single ground-truth capability, namely, computing the GCD. Yet, the theoretical portion of our work holds for any ground-truth $F^{\ast}$ that admits an Interactive Proof system. Training large Self-Proving models for more complex ground-truths will likely pose additional practical learning challenges. With that said, we stress that generating accepting transcripts for use in Transcript Learning is distinct from these learning challenges. Collecting accepting transcripts is a purely computational task, and can even be done “offline” prior to the model’s training.
Additionally, in our current learning methods, each individual ground-truth capability requires training a separate Self-Proving model. It would be interesting to adapt our definition and methods to deal with a single generalist Self-Proving model that proves its correctness to multiple verifiers of different ground-truths.
7 Conclusions
Trust between a learned model and its user is fundamental. In recent decades, Interactive Proofs (Goldwasser et al., 1985) have emerged as a general theory of trust established via verification algorithms. This work demonstrates that models can learn to formally prove their answers in an Interactive Proof system. We call models that possess this capability Self-Proving.
The definition of Self-Proving models forms a bridge between the rich theory of Interactive Proofs and the contemporary topic of Trustworthy ML. Interactive Proofs offer formal worst-case soundness guarantees; thus, users of Self-Proving models can be confident when their models generate correct answers—and detect incorrect answers with high probability.
We demonstrate the theoretical viability of our definition with two generic learning algorithms: Transcript Learning (TL) and Reinforcement Learning from Verifier Feedback (RLVF). The analyses of these algorithms is informed by techniques from theories of learning, RL, and computational complexity. This work can be extended in several directions: finding conditions for the convergence of RLVF, improving sample complexity bounds for TL, or designing altogether different learning algorithms (for example, by taking advantage of properties of the verifier).
To better understand the training dynamics of (Annotated) TL, we train Self-Proving transformers for the Greatest Common Divisor (GCD) problem. We train a small (6.3M parameter) transformer that learns to generate correct answers and proofs with high accuracy. Facing forward, we note that Interactive Proofs exist for capabilities far more complex than the GCD (Shamir, 1992); scaling up our experiments is the next step towards bringing Self-Proving models from theory to practice.
Acknowledgments
We are grateful to Micah Carroll, Mojtaba Yaghoobzadeh, Avishay Tal, and anonymous reviewers for their helpful comments. This research was supported by DARPA-TA1 under grant no. HR001119S0076, and by the Simons Collaboration on the Theory of Algorithmic Fairness.
Appendix A A definition for general loss functions and one-to-many relations
We present a variant of Self-Proving models (Definition 3.4) generalized in two ways.
@@ Appendix A A definition for general loss functions and one-to-many relations
In Definition 3.1 we implicitly use the 0-1 loss when measuring the correctness of a model: For any $x\in X$, we measure only whether the model generated the correct output $y=F^{\ast}(x)$, but not how “far” the generated $y$ was from $F^{\ast}(x)$. It is often the case in machine learning that we would be satisfied with models that generate a “nearly-correct” output. This is formalized by specifying a loss function $\ell\colon\Sigma^{\ast}\times\Sigma^{\ast}\to[0,1]$ and measuring the probability that $\ell(x,y)$ is smaller than some threshold $\lambda\in[0,1)$, where $x$ is drawn from the input distribution $\mu$, and $y$ is generated by the model when given input $x$.
@@
In the notation of Section 3, the prime-counting task would be captured by the ground-truth function
@@
Per Definition 3.1, any output other than $F^{\ast}(x)$ is “just as incorrect” as any other. Yet, we might prefer outputs that are closer to the correct answer, say, in $L_{1}$ norm. This preference can be captured by the following bounded loss function
@@
More generally, we relax Definition 3.1 to capture approximate correctness as follows.
Definition A.1 (Approximate correctness).
@@
In Section 3, we focused on the setting of models of a ground-truth function $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$. That is, when each input $x$ has exactly one correct output, namely $F^{\ast}(x)$. A more general setting would be to consider a ground-truth relation $L\subseteq\Sigma^{\ast}\times\Sigma^{\ast}$. Then, we say that $y$ is a correct output for $x$ if $(x,y)\in L$. Importantly, this allows a single $x$ to have many possible correct outputs, or none at all.
Note that we must take care to choose a loss function $\ell$ that captures correctness with respect to the relation $L$, i.e., $\ell(x,y)=0$ if and only if $(x,y)\in L$. Equivalently, any loss function $\ell$ induces a relation $L\coloneqq\{(x,y)\ |\ \ell(x,y)=0\}$. Therefore, our relaxation to approximate-correctness Definition A.1 already captures the setting of one-to-many relations, since an input $x$ may have multiple $y^{\ast}$ such that $\ell(x,y^{\ast})=0$.
A.1 The general definition

### v4

@@ 3.2 Self-Proving Models
We define the Verifiability of a model $P_{\theta}$ with respect to an input distribution $\mu$ and a verifier $V$. Intuitively, Verifiability captures the ability of the model to prove the correctness of its answer $y\sim P_{\theta}(x)$, when the input $x$ is sampled from $\mu$. We refer to models capable of proving their own correctness as Self-Proving models. Notice that, as in Definition˜3.2, the verifier is fixed and agnostic to the choice of the Self-Proving model.
@@
Fix a verifier $V$ for a ground-truth $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$ as in Definition˜3.2, and a distribution $\mu$ over inputs $\Sigma^{\ast}$. The Verifiability of a model $P_{\theta}\colon\Sigma^{\ast}\to\Sigma^{\ast}$ is defined as
@@
Notice that the ground-truth $F^{\ast}$ does not appear in Definition˜3.4 except for the first sentence. Indeed, once it is established that $V$ is a verifier for $F^{\ast}$ (as per Definition˜3.2), then Verifiability w.r.t $V$ implies correctness w.r.t $F^{\ast}$: Consider any input distribution $\mu$, ground-truth $F^{\ast}$, and a verifier $V$ for $F^{\ast}$ with soundness error $s$. By a union bound, if a model $P_{\theta}$ is $\beta$-Verifiable, then it is $(\beta-s)$-correct. That is to say, Verifiability is formally a stronger guarantee than correctness when $V$ has small soundness error $s$.
As depicted in Figure˜1, a Self-Proving model $P_{\theta}$ plays a dual role: first, it generates an output $y\sim P_{\theta}(x)$, and then it proves the correctness of this output to $V$. Note also that Verifiability is a feature of a model, unlike completeness and soundness which are features of a verifier (see Table˜1).
@@
3.3 A Definition for General Loss Functions and One-To-Many Relations
We present variants of Self-Proving models (Definition˜3.4) generalized to one-to-many relations, and general bounded loss functions. While these generalizations provide a richer framework that may accommodate a wider range of applications, the theorems in this paper are based on the forgoing Definition˜3.4, which captures the essential properties while remaining mathematically manageable.
@@ 3.3 A Definition for General Loss Functions and One-To-Many Relations
In Definition˜3.1 we implicitly use the 0-1 loss when measuring the correctness of a model: For any $x\in X$, we measure only whether the model generated the correct output $y=F^{\ast}(x)$, but not how “far” the generated $y$ was from $F^{\ast}(x)$. It is often the case in machine learning that we would be satisfied with models that generate a “nearly-correct” output. This is formalized by specifying a loss function $\ell\colon\Sigma^{\ast}\times\Sigma^{\ast}\to[0,1]$ and measuring the probability that $\ell(x,y)$ is smaller than some threshold $\lambda\in[0,1)$, where $x$ is drawn from the input distribution $\mu$, and $y$ is generated by the model when given input $x$.
@@
In the notation of Section˜3, the prime-counting task would be captured by the ground-truth function
@@
Per Definition˜3.1, any output other than $F^{\ast}(x)$ is “just as incorrect” as any other. Yet, we might prefer outputs that are closer to the correct answer, say, in $L_{1}$ norm. This preference can be captured by the following bounded loss function
@@
More generally, we relax Definition˜3.1 to capture approximate correctness as follows.
Definition 3.6 (Approximate correctness).
@@
In Section˜3, we focused on the setting of models of a ground-truth function $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$. That is, when each input $x$ has exactly one correct output, namely $F^{\ast}(x)$. A more general setting would be to consider a ground-truth relation $L\subseteq\Sigma^{\ast}\times\Sigma^{\ast}$. Then, we say that $y$ is a correct output for $x$ if $(x,y)\in L$. Importantly, this allows a single $x$ to have many possible correct outputs, or none at all.
Note that we must take care to choose a loss function $\ell$ that captures correctness with respect to the relation $L$, i.e., $\ell(x,y)=0$ if and only if $(x,y)\in L$. Equivalently, any loss function $\ell$ induces a relation $L\coloneqq\{(x,y)\ |\ \ell(x,y)=0\}$. Therefore, our relaxation to approximate-correctness Definition˜3.6 already captures the setting of one-to-many relations, since an input $x$ may have multiple $y^{\ast}$ such that $\ell(x,y^{\ast})=0$.
3.3.1 The General Definition